# Adaptive Debiased SGD in High-dimensional Generalized Linear Models with Streaming Data

### v2

@@ 3.2 Asymptotic normality
By Theorems 3.1 and 3.2, we can establish the asymptotic normality of ${\widehat{\bm{\beta}}}^{(n)}_{j,\text{de}}$ defined in (13).
@@
With the choice of hyper-parameters in (20) and (21), given any fixed $j\in[p],$
@@
Compared to existing works in high-dimensional inferences, our theoretical results are different in two aspects. On one hand, the conditions in Theorem 3.3 are made at the population level. In contrast, most of the existing inference methods (van de Geer et al.,, 2014; Ning and Liu,, 2017; Ma et al.,, 2021; Shi et al.,, 2021; Luo et al., 2023a, ) rely on the bounded individual probability condition. For example, in logistic regression, the bounded individual probability condition takes the form that $\mathbb{P}(y_{i}=1|\bm{x}_{i})\in(c,1-c)$ for $i\in[n]$ and some positive universal constant $c$. However, such a condition can be easily violated if $\bm{x}_{i}$ follows a normal distribution with a slightly large variance. To the best of our knowledge, except for Guo et al., (2021) and Cai et al., (2023), most existing methods require the bounded individual probability condition. Without the bounded individual probability condition, our method is shown to provide valid inference under the Gaussian design. The removal of such a condition is not only a technical advancement, but also a methodological improvement as our method is able to handle more general cases.
On the other hand, we require slightly stronger conditions in terms of the sparsity levels $s_{0}$ and $s_{1}$, sample size $n$, and the dimension of features $p$. This is the trade-off between online and offline algorithms. Specifically, the online estimator may not be the optimizer of some loss functions. Hence, some optimality conditions, such as the Karush–Kuhn–Tucker (KKT) condition, cannot be used. Therefore, we need to impose slightly stronger conditions compared to its offline counterpart so as to establish the asymptotic properties; see Theorem 4.10 in Deshpande et al., (2023) and Theorem 1 in Han et al., (2024).
@@ 4 Numerical studies
In this section, we conduct simulation studies to examine the finite-sample performance of the ADL in high-dimensional logistic regression. Besides that, we provide a real data example to show the ability of our method in dealing with ultra-high dimensional online data. All experiments are performed with an 11th Gen Intel Core i7-11370H processor and 16GB memory storage.
4.1 Simulated data
@@
We randomly generate a sequence of $n$ i.i.d. copies $\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$ realized from the logistic regression model $\mathbb{P}(y=1\lvert\bm{x};\bm{\beta}^{*})=1-\mathbb{P}(y=0\lvert\bm{x};\bm{\beta}^{*})={\exp(\bm{x}^{\intercal}\bm{\beta}^{*})}/\{1+\exp(\bm{x}^{\intercal}\bm{\beta}^{*})\}$, where $\bm{x}\sim\mathcal{N}(0,\bm{\Sigma})$. Recall that the cardinality of the support set of $\bm{\beta}^{*}$ is $s_{0}$. We divide the support set into two halves $\{D_{1},D_{2}\}$, with true coefficients $\bm{\beta}^{*}_{D_{1}}=1$ and $\bm{\beta}^{*}_{D_{2}}=-1$ respectively. Both $D_{1}$ and $D_{2}$ are randomly selected. We conduct two experiments: one is the comparison between the ADL and some existing methods to show the advantage of the ADL in terms of time complexity. Another is used to indicate the low space complexity of the ADL.
@@
In our first experiment, we choose $n=200,p=500$ and $s_{0}=6$ and compare the performance of the ADL with offline debiased lasso (deLasso) in van de Geer et al., (2014), link-specific weighting (LSW) method proposed in Cai et al., (2023) and online debiased lasso (ODL) proposed in Luo et al., 2023a . In addition, we check the robustness of these methods with respect to violating the bounded individual probability condition or not. For this reason, we consider two types of covariance matrix $\bm{\Sigma}=0.1\times\{0.5^{|i-j|}\}_{i,j=1,\dots,p}$ and $\bm{\Sigma}=\{0.5^{|i-j|}\}_{i,j=1,\dots,p}$. This setting is similar to the setting in Cai et al., (2023). The former will ensure the bounded individual probability condition holds, while the latter will break this condition since the resulting probability can be pretty close to 00. Besides that, the consideration of other covariance matrices, such as the block-wise Toeplitz matrix, as well as the choice of the hyper-parameters is given in the supplementary material.
@@
The performance is based on four criteria: (1) Absolute bias: the absolute distance between estimator and ground truth. (2) Converge probability of 95% confidence interval (3) Length of 95% confidence interval and (4) the computation times. There are three different categories among $\bm{\beta}^{*}_{i},i\in[p]$, that is $0,1$ and $-1$. We choose three elements from each category and calculate the corresponding debiased estimator respectively. The averaged performances for each category are recorded. The experiment is conducted with 200 replications and the results are summarized in Tables 2 and 3.
@@
Table 2 presents the results when the covariance matrix $\bm{\Sigma}=0.1\times\{0.5^{|i-j|}\}_{i,j=1,\dots,p}$. Notably, all four methods are able to achieve the desired coverage probability. Our ADL estimator takes the minimum time and can still achieve the comparable length of the confidence interval. In Table 3, when the bounded individual probability condition is violated, both deLasso and its online variant ODL fail to achieve $95\%$ coverage probability. This phenomenon has also been observed in Cai et al., (2023). In contrast, LSW and ADL are still able to achieve $95\%$ coverage probability. This result further supports our finding that similar to LSW, the ADL does not depend on the bounded individual probability condition. Compared with LSW, the ADL method is around 50 to 60 times faster with similar statistical efficiency. In summary, the ADL enjoys a small time complexity and still yields high statistical accuracy.
@@
In the second setting, we aim to demonstrate the low space complexity of the ADL. We increase the value of $p$ to $20000$ and choose $n=1000,s_{0}=20$ with $\bm{\Sigma}=\{0.5^{|i-j|}\}_{i,j=1,\dots,p}$. Under this circumstance, the computation of deLasso, LSW, or ODL is out of memory and only ADL is computable. Similar to the first setting, we construct the confidence intervals for three randomly selected parameters from each category of $\bm{\beta}^{*}$. The averaged result in each category is summarized in Figure 3. In Figure 3, the lines and shadow stand for the change of debiased estimator and the length of 95% confidence interval along with the increase of data size respectively. We can observe that the estimator is closed to the corresponding true value in every category, which shows the accuracy of our method. Another unexpected observation is that the length of the confidence interval decreases as the number of data increases. The coverage probability also achieves 95% for each case and the detailed results can be found in the supplementary material. The second experiment shows the ADL has low space complexity but still maintains the desired statistical accuracy.
@@
Table 2: Simulation results over 200 replications with $n=200$, $p=500$, $s_{0}=6$, $\bm{\Sigma}=0.1\times\{0.5^{|i-j|}\}_{i,j=1,\dots,p}$.
@@
Absolute bias 0 0.410 0.368 0.066 0.043 0.038 0.716 0.510 0.435
1 0.436 0.532 0.259 0.219 0.181 0.707 0.486 0.417
-1 0.411 0.521 0.231 0.204 0.178 0.707 0.545 0.436
Coverage probability 0 0.953 0.988 0.971 0.963 0.958 0.968 0.963 0.950
1 0.930 0.925 0.963 0.951 0.946 0.963 0.975 0.961
-1 0.952 0.945 0.975 0.958 0.953 0.978 0.956 0.950
Confidence interval length 0 2.036 2.260 4.061 2.297 1.771 3.903 2.665 2.143
1 2.036 2.265 4.059 2.293 1.766 3.916 2.665 2.151
-1 2.045 2.284 4.09 2.304 1.776 3.855 2.654 2.140
Time $(s)$ 7.214 20.22 0.539 0.324
Table 3: Simulation results averaged over 200 replications with $n=200$, $p=500$, $s_{0}=6$, $\bm{\Sigma}=\{0.5^{|i-j|}\}_{i,j=1,\dots,p}$.
@@
Absolute bias 0 0.114 0.113 0.018 0.015 0.013 0.509 0.344 0.295
1 0.325 0.260 0.556 0.505 0.488 0.520 0.369 0.302
-1 0.326 0.259 0.574 0.522 0.481 0.539 0.384 0.316
Coverage probability 0 0.983 0.996 1.000 0.999 0.997 0.948 0.965 0.956
1 0.598 0.936 0.956 0.790 0.646 0.928 0.940 0.953
-1 0.610 0.946 0.958 0.828 0.676 0.918 0.926 0.928
Confidence interval length 0 0.753 1.207 1.834 1.354 1.133 2.487 1.751 1.436
1 0.769 1.261 1.833 1.355 1.131 2.546 1.784 1.458
-1 0.767 1.271 1.828 1.348 1.130 2.496 1.759 1.449
Time $(s)$ 6.550 20.04 0.628 0.365
@@
Figure 3: Simulation results averaged over 200 replications (each takes around 53.54s) with $n=1000$, $p=20000$, $s_{0}=20$, $\bm{\Sigma}=\{0.5^{|i-j|}\}_{i,j=1,\dots,p}$.
4.2 Real data example: Identifying malicious URLs
With the rapid development of online technology, it has become more convenient for people to communicate with each other. Nowadays, human communication is not only limited to emails but also through websites and social networking platforms. However, this has also created opportunities for criminals to take advantage of the situation. Every day we still receive lots of spam emails, malicious links, and phishing websites from the Internet. How to identify fake information is important, which can help us to avoid the leakage of personal information and the loss of property. For this purpose, Ma et al., (2009) considered an online algorithm to predict malicious URLs. The data is collected in 120 days, resulting in 2.4 million data and around 3.2 million features. The authors compared several methods and reported the cumulative prediction errors. Specifically, they considered the logistics regression with the stochastic gradient descent, whose prediction error is below 5%.
Different from the prediction error, we are more interested in which attributes are significant to make the malicious URL distinguished from the benign one. We take one further step to conduct inference with the ADL method. We evaluate our algorithm in two aspects: accuracy and computational efficiency. However, it is pretty hard to judge accuracy since all features are anonymized. For this reason, we add one fake feature, which is randomly generated from the standard normal distribution, as a benchmark. In addition to the fake feature, we also conduct inference for features 16 and 32. The results of the first-day data (with the size 20,000) are presented in Figure 4.
Similar to Figure 3, the lines and shadow show the ADL estimator and the 95% confidence interval respectively. According to the right panel, we observe that the confidence interval of the fake feature always contains zero, which indicates the accuracy of ADL. Besides that, the results from the left and middle panels imply that Feature 16 is significant while Feature 32 is not under the significance level 0.05.
In terms of computational efficiency, the online algorithm should be able to give the response for the current data prior to the arrival of the next data. The data in Ma et al., (2009) is collected in an online fashion, with approximately $4.32s$ for each. In contrast, the ADL costs $0.3s$ for processing each single data. This result reveals the efficiency of the ADL and its ability to conduct online inference with ultra-high dimensional data.
Figure 4: 95% confidence interval estimate for Feature 16 (left), Feature 32 (middle), and Fake Feature (right).

### v3

@@ 3.2 Asymptotic normality
By Theorems 3.1 and 3.2, we establish the asymptotic normality of ${\widehat{\bm{\beta}}}^{(n)}_{j,\text{de}}$ defined in (13).
@@
With the choice of hyper-parameters according to (S.1)–(S.6) in the supplementary material, given any fixed $j\in[p],$
@@
Compared to existing work in high-dimensional inference, our theoretical results are different in two aspects. On one hand, the assumptions in Theorem 3.3 are imposed at the population level. In contrast, most of the existing inference methods (van de Geer et al.,, 2014; Ning and Liu,, 2017; Ma et al.,, 2021; Shi et al.,, 2021; Luo et al., 2023a, ) rely on the bounded individual probability condition. For example, in logistic regression, the bounded individual probability condition takes the form $\mathbb{P}(y_{i}=1|\bm{x}_{i})\in(c,1-c)$ for $i\in[n]$ and some positive universal constant $c$. This difference in assumptions is primarily due to stochastic optimization in our work, where the objective function is defined at the population level. Consequently, we do not rely on assumptions at the sample level. When $\bm{x}_{i},i\in[n]$ follows a normal distribution with a slightly larger variance, the bounded individual probability condition may not be satisfied. To the best of our knowledge, except for Guo et al., (2021) and Cai et al., (2023), most existing methods require the bounded individual probability condition. Therefore, both our method and the one proposed in Cai et al., (2023) can still provide valid inference under Gaussian design, as evidenced in our numerical studies.
On the other hand, we require slightly stronger conditions in terms of the sparsity levels $s_{0}$ and $s_{1}$, sample size $n$, and the dimension of features $p$. This is the trade-off between online and offline algorithms. Specifically, the online estimator may not be the optimizer of some objective functions. Hence, some optimality conditions, such as the Karush–Kuhn–Tucker (KKT) condition, cannot be used. Therefore, we need to impose slightly stronger conditions compared to its offline counterpart so as to establish the asymptotic properties. Similar consideration can be also found in the existing theoretical results related to online high-dimensional inference, such as Theorem 4.10 in Deshpande et al., (2023) and Theorem 1 in Han et al., (2024).
@@ 4 Simulation studies
In this section, we conduct simulation studies to examine the finite-sample performance of the ADL in high-dimensional logistic regression. All experiments, including the real data example in Section 5, are performed with an 11th Gen Intel Core i7-11370H processor and 16GB memory storage.
@@
Simulation settings: We randomly generate a sequence of $n$ independent and identically distributed copies $\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$ from the logistic regression model $\mathbb{P}(y=1\lvert\bm{x};\bm{\beta}^{*})=1-\mathbb{P}(y=0\lvert\bm{x};\bm{\beta}^{*})={\exp(\bm{x}^{\intercal}\bm{\beta}^{*})}/\{1+\exp(\bm{x}^{\intercal}\bm{\beta}^{*})\}$, where $\bm{x}\sim\mathcal{N}(0,\bm{\Sigma})$. Recall that the cardinality of the support set of $\bm{\beta}^{*}$ is $s_{0}$. We divide the support set into two halves $\{S_{1},S_{2}\}$, with true coefficients $\bm{\beta}^{*}_{S_{1}}=1$ and $\bm{\beta}^{*}_{S_{2}}=-1$ respectively. Both $S_{1}$ and $S_{2}$ are randomly selected. We conduct two experiments: one is the comparison between the ADL and some existing methods to show the advantage of the ADL in terms of time complexity. Another is used to indicate the low space complexity of the ADL.
@@
In our first experiment, we choose $n=200,p=500$ and $s_{0}=6$. We compare the performance of the ADL with offline debiased lasso (deLasso) in van de Geer et al., (2014), link-specific weighting (LSW) method proposed in Cai et al., (2023) and online debiased lasso (ODL) proposed in Luo et al., 2023a . In addition, we check the robustness of these methods with respect to violating the bounded individual probability condition or not. For this reason, we consider two types of covariance matrix $\bm{\Sigma}=0.1\times\{0.5^{|i-j|}\}_{i,j=1,\dots,p}$ and $\bm{\Sigma}=\{0.5^{|i-j|}\}_{i,j=1,\dots,p}$. This setting is similar to the setting in Cai et al., (2023). The former will ensure the bounded individual probability condition holds, while the latter will break this condition since the resulting probability can be pretty close to 00.
@@
The performance is based on four criteria: (1) Absolute bias: the absolute distance between estimator and ground truth. (2) Coverage probability of 95% confidence interval (3) Length of 95% confidence interval and (4) the computation times. There are three different categories among $\bm{\beta}^{*}_{i},i\in[p]$, that is $0,1$ and $-1$. We choose three elements from each category and calculate the corresponding debiased estimator respectively. The averaged performances for each category are recorded. The experiment is conducted with 500 replications and the results are summarized in Tables 2 and 3.
@@
Table 2 presents the results when the covariance matrix $\bm{\Sigma}=0.1\times\{0.5^{|i-j|}\}_{i,j=1,\dots,p}$. Notably, all four methods are able to achieve the desired coverage probability. Our ADL estimator takes the minimum time and can still achieve the comparable length of the confidence interval. In Table 3, when the bounded individual probability condition is violated, both deLasso and its online variant ODL fail to achieve $95\%$ coverage probability. This phenomenon has also been observed in Cai et al., (2023). In contrast, LSW and ADL are still able to achieve $95\%$ coverage probability. This result further supports our finding that similar to LSW, the ADL does not depend on the bounded individual probability condition. Compared with LSW, the ADL estimator is around 60 to 70 times faster with similar statistical efficiency. In summary, the ADL enjoys a small time complexity and still yields high statistical accuracy.
@@
In the second experiment, we aim to demonstrate the low space complexity of the ADL. We increase the value of $p$ to $20000$ and choose $n=1000,s_{0}=20$ with $\bm{\Sigma}=\{0.5^{|i-j|}\}_{i,j=1,\dots,p}$. Under this circumstance, the computation of deLasso, LSW, or ODL is out of memory and only ADL is computable. Similar to the first setting, we construct the confidence intervals for three randomly selected parameters from each category of $\bm{\beta}^{*}$. The averaged result in each category is summarized in Figure 3. In Figure 3, the lines and shadow stand for the change of debiased estimator and the length of 95% confidence interval along with the increase of data size respectively. We observe that the estimator is close to the corresponding true value in every category, which shows the accuracy of our method. In addition, the length of the confidence interval decreases as the sample size increases, and the coverage probability is close to 95% in each case, with detailed results provided in the supplementary material.
The second experiment indicates that the ADL has low space complexity while maintaining the desired statistical accuracy.
@@
Table 2: Simulation results over 500 replications with $n=200$, $p=500$, $s_{0}=6$, $\bm{\Sigma}=0.1\times\{0.5^{|i-j|}\}_{i,j=1,\dots,p}$.
@@
Absolute bias 0 0.409 0.374 0.032 0.029 0.026 1.360 0.609 0.486
1 0.397 0.485 0.228 0.180 0.165 1.326 0.638 0.497
-1 0.414 0.511 0.246 0.206 0.189 1.308 0.611 0.484
Coverage probability 0 0.952 0.972 0.980 0.971 0.966 0.945 0.974 0.960
1 0.962 0.936 0.978 0.962 0.962 0.969 0.969 0.958
-1 0.955 0.930 0.971 0.958 0.957 0.961 0.967 0.965
Confidence interval length 0 2.032 2.200 3.076 2.333 1.970 6.503 3.326 2.527
1 2.034 2.197 3.074 2.337 1.972 6.461 3.334 2.539
-1 2.032 2.202 3.070 2.333 1.969 6.470 3.325 2.528
Time $(s)$ 7.954 21.017 1.228 0.310
Table 3: Simulation results averaged over 500 replications with $n=200$, $p=500$, $s_{0}=6$, $\bm{\Sigma}=\{0.5^{|i-j|}\}_{i,j=1,\dots,p}$.
@@
Absolute bias 0 0.115 0.117 0.013 0.011 0.010 0.830 0.364 0.298
1 0.337 0.260 0.583 0.515 0.486 0.811 0.399 0.311
-1 0.338 0.256 0.574 0.511 0.481 0.774 0.395 0.310
Coverage probability 0 0.982 0.995 0.999 0.999 0.997 0.958 0.973 0.960
1 0.578 0.928 0.948 0.793 0.672 0.952 0.945 0.940
-1 0.573 0.940 0.947 0.802 0.666 0.962 0.951 0.944
Confidence interval length 0 0.749 1.206 1.836 1.353 1.130 3.868 1.953 1.492
1 0.763 1.243 1.827 1.351 1.128 3.933 1.978 1.511
-1 0.763 1.245 1.833 1.347 1.129 3.890 1.967 1.502
Time $(s)$ 8.127 19.744 1.189 0.293
@@
Figure 3: Simulation results averaged over 500 replications (each takes around 42.452s) with $n=1000$, $p=20000$, $s_{0}=20$, $\bm{\Sigma}=\{0.5^{|i-j|}\}_{i,j=1,\dots,p}$.
5 Real data example: spam email classification
With the advancement of digital communication, email has become an indispensable tool for both personal and professional correspondence. However, this widespread adoption has also attracted malicious actors who exploit email systems to distribute spam, phishing attempts, and fraudulent content. Identifying spam emails and understanding the features that characterize them are crucial for protecting users from potential security threats, financial fraud, and privacy breaches.
For this purpose, we analyze a dataset on spam email classification [^1] https://www.kaggle.com/datasets/purusinghvi/email-spam-classification-dataset. This dataset consists of 83,446 email messages with binary labels for spam or non-spam. We construct features by counting the frequencies of unigrams and bigrams from the email content and then applying a log transformation. In addition, we partition the data into training and test sets to evaluate our algorithm’s prediction accuracy.
Specifically, we split the data into 80% for training and 20% for testing. As a result, the feature dimension $p\approx 3.6$ million, and the training data size $n\approx 66,000$. We then apply our method to the training set and evaluate the prediction accuracy on the test set using the estimator derived from the training step. It is important to note that this dataset is ultra-high dimensional and the existing offline methods are computationally prohibitive, so we only show results from our proposed ADL method.
We select the terms of our interests: “investment”, “schedule”, and “per cent” to conduct statistical inference, with their confidence bands shown in the upper left, upper right, and bottom left of Figure 4. For each term, the corresponding confidence interval gets narrower as more data are processed. Notably, the term “investment” consistently shows a positive coefficient, with confidence intervals above zero after processing 950 data points. In contrast, the term “schedule” presents a significantly negative relationship, with its confidence intervals falling below zero after 1100 data points. These findings align with our intuition: emails containing “investment” are more likely to be financial scams, while “schedule” typically appears in trusted business communications, like scheduling meetings. Unlike “investment” and “schedule,” the term “per cent” shows a neutral influence, with confidence bands covering zero. This phenomenon suggests that “per cent” is a common term in both spam and non-spam emails. Thus, it has a limited ability to distinguish spam from non-spam emails. The bottom-right panel of Figure 4 shows the prediction error rate evaluated on the test set as training sample size increases. As the ADL estimator is updated alongside the increasing training sample, the prediction error rate steadily decreases from approximately 0.47 to 0.15. This result demonstrates the effectiveness of our method from the perspective of prediction accuracy.
Figure 4: 99% confidence interval estimate for “investment”, “schedule”, and “per cent” (upper left, upper right, and bottom left), where lines and shaded areas represent the traces of the ADL estimator and 99% confidence interval respectively. The prediction error evaluated on the test set, as training sample size increases, is shown in the bottom right.